# On the dynamics of leapfrogging vortex rings

### v1

@@ 4.3. Turkington’s-type Lemma
$-\frac{25C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|\leqslant\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t).$
@@
00 0 $\leqslant\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)+\frac{25C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|$
+$\leqslant\int_{{\mathbb{R}}^{2}_{+}}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{25C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|+\mathcal{O}(1),$
@@
+$\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}\psi^{\varepsilon}(t,x)+\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\\ \leqslant\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{26C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|.$
@@
$\leqslant$ $\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{26C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|$
@@
$+\frac{26C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|$
@@
$+\frac{26C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|$
$\leqslant$ $\frac{\gamma^{2}}{\pi}h_{0}(5+26C^{\prime}_{F}T_{0})\ln|\ln\varepsilon|,$
where we used Lemma 4.6 and Lemma 4.4. Choosing $A=\frac{1}{\pi}\gamma r^{*}(5+26C^{\prime}_{F}T_{0})$ gives the first estimate in the lemma for any $\varepsilon\in(0,\varepsilon_{0}]$.
@@
where we have used $\sqrt{x_{r}y_{r}}-b_{r}^{\varepsilon}(t)=\mathcal{O}(1/\sqrt{|\ln\varepsilon|})$ for any $x,y\in\operatorname{supp}\omega^{\varepsilon}(t)$ and (4.3). Therefore
@@ 4.4. Bootstrap argument and ends of the proof of the main results of this section
$\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x\leqslant\frac{1}{2}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
@@
By Lemma 4.11 applied to $\rho(\varepsilon)=\frac{B}{|\ln\varepsilon|-2\ln|\ln\varepsilon|}$, there exists a set $\widetilde{\Omega}_{i}^{\varepsilon}(t)$ such that
@@
$\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\leqslant\frac{h_{0}}{r^{*}}\rho(\varepsilon)\ln|\ln\varepsilon|=\frac{h_{0}}{r^{*}}\frac{B\ln|\ln\varepsilon|}{|\ln\varepsilon|-2\ln|\ln\varepsilon|}\leqslant\frac{1}{2}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|},$
@@
(4.7) $\frac{Bh_{0}}{r^{*}}=20(5+26C_{F}^{\prime}T_{0})\frac{h_{0}}{r*}\leqslant\frac{1}{2}.$
The above condition is the constraint about $h_{0}$ and $T_{0}$ in the main results stated at the beginning of this section. Assuming for instance $T_{0}\leqslant 1/(26C_{F}^{\prime})$ and $h_{0}\leqslant r^{*}/240$, all the results of this section reads as: there exists $\varepsilon_{0}>0$ depending only on $h_{0}$ such that for any $\varepsilon\in(0,\varepsilon_{0}]$ and every $i\in\{1,\ldots,N\}$, all the inequalities stated hold true for all $t\in[0,T_{\varepsilon}]$. In particular, Proposition 4.1 and respectively Lemma 4.2 come directly from Lemma 4.11 and resp. Lemma 4.4. Finally, Corollary 4.3 is a direct consequence of Lemma 4.4 combined with the definition of $\Omega_{i}^{\varepsilon}$ (4.2).
@@
The constant $h_{0}$ was chosen small enough so that some constants are absorbed, see (4.7), which explained why we have tracked in all this section the dependence on $h_{0}$. The possibility to choose $h_{0}$ small will not be used anymore, therefore, $h_{0}$ is fixed as $h_{0}=\min(r^{*}/240,1/2)$ and we will not specify the dependence of the constants on $h_{0}$ in the sequel.
@@
which means that most of the mass can be found in an area of radius $\varepsilon^{1-\eta}$. However, the estimate on the quantity of outside mass being $\mathcal{O}\left(\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)$, with the localization assumptions we can only prove that
@@
and not better. Therefore, the center of mass can be quite far from where the filament is, compared to the size of the core of the filament, due to the possible filamentation of the filament.
@@ 5. Equation of the motion
In Lemma 4.9 together with the proof that $\widetilde{T}_{\varepsilon}=T_{\varepsilon}$ (see the end of the previous section), we have obtained the equation of the motion of the radial component of $b^{\varepsilon}_{i}(t)$. In this section, we now obtain the full equation for both components, then make the link with $(X_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$.
@@ 5.1. Motion of the center of mass
Keeping for now the notations of a single filament in an exterior field, we compute the following equation for the motion of the center of mass of the vortex ring.
@@
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$, there exists $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$\left|\frac{\mathrm{d}}{\mathrm{d}t}b_{i}^{\varepsilon}(t)-\frac{1}{|\ln\varepsilon|}F_{i}^{\varepsilon}(t,b_{i}^{\varepsilon}(t))-\gamma\frac{e_{z}}{4\pi b_{i,r}^{\varepsilon}(t)}\right|\leqslant C(1+T_{0})\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
@@
Recall that
@@
$\hskip 28.45274pt+\frac{1}{\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}u_{L}^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\hskip 56.9055pt+\frac{1}{\gamma|\ln\varepsilon|}\int u_{R}^{\varepsilon}(t,x)x_{r}\omega(t,x)\mathrm{d}x$
$\hskip 85.35826pt+\frac{1}{\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}F^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$:=A_{1}+A_{2}+A_{3}+A_{4},$
@@
$A_{4}=\frac{1}{|\ln\varepsilon|}F^{\varepsilon}(t,b^{\varepsilon}(t))+\frac{1}{\gamma|\ln\varepsilon|}\int\big{(}F^{\varepsilon}(t,x)-F(t,b^{\varepsilon}(t))\big{)}x_{r}\omega(t,x)\mathrm{d}x$
then using from Lemma 3.1 that $F$ is Lipschitz and Lemma 4.2
@@
$\hskip 28.45274pt-\frac{e_{z}}{2\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}\left(\frac{1}{x_{r}^{2}}-\frac{1}{|b_{r}^{\varepsilon}(t)|^{2}}\right)\psi^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x.$
@@
then by using Corollary 4.10 we obtain that
$\left|A_{2}-\gamma\frac{e_{z}}{4\pi b_{r}^{\varepsilon}(t)}\right|\leqslant C(1+T_{0})\frac{\ln|\ln\varepsilon|}{\ln\varepsilon}.$
Finally, we show that the velocity created by the other vortices $F_{i}^{\varepsilon}$ behaves as the point vortex interaction in the classical point vortex system (see [40, 41]).
@@
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$, there exists $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have
@@
$\hskip 28.45274pt-\frac{e_{z}}{4\pi b_{i,r}^{\varepsilon}(t)}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\sqrt{\frac{y_{r}}{b_{i,r}^{\varepsilon}(t)}}\ln|b_{i}^{\varepsilon}(t)-y|y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y$
$\hskip 56.9055pt+\frac{1}{b_{i,r}^{\varepsilon}(t)}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\nabla^{\perp}_{x}S(b_{i}^{\varepsilon}(t),y)y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y$
$:=A_{1}+A_{2}+A_{3}.$
@@
5.2. Limit trajectories
We prove in this section that $(b_{i}^{\varepsilon})_{i}$, defined in (1.12), are located close to $(X_{i}^{\varepsilon})_{i}$ and $(\widetilde{X}_{i}^{\varepsilon})_{i}$, defined in (1.8) and resp. in (1.11).
First, we prove that $(X_{i}^{\varepsilon})_{i}$ and $(\widetilde{X}_{i}^{\varepsilon})_{i}$ are located in $\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}(d_{0},h_{0})$, see (3.4) for the definition of this set, and remains close.

### v2

@@ 4.3. Turkington’s-type Lemma
$-\frac{(24T_{0}C^{\prime}_{F}+\frac{1}{4})\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|\leqslant\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t).$
@@
00 0 $\leqslant\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)+\frac{(24T_{0}C^{\prime}_{F}+\frac{1}{4})\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|$
+$\leqslant\int_{{\mathbb{R}}^{2}_{+}}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{(24T_{0}C^{\prime}_{F}+\frac{1}{4})\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|+\mathcal{O}(1),$
@@
+$\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}\psi^{\varepsilon}(t,x)+\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\\ \leqslant\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{(24T_{0}C^{\prime}_{F}+\frac{1}{2})\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|.$
@@
$\leqslant$ $\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{(24T_{0}C^{\prime}_{F}+\frac{1}{2})\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|$
@@
$+\frac{(24T_{0}C^{\prime}_{F}+\frac{1}{2})\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|$
@@
$+\frac{(24T_{0}C^{\prime}_{F}+\frac{1}{2})\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|$
$\leqslant$ $\frac{\gamma^{2}}{\pi}h_{0}(5+24C^{\prime}_{F}T_{0})\ln|\ln\varepsilon|,$
where we used Lemma 4.6 and Lemma 4.4. Choosing $A=\frac{1}{\pi}\gamma r^{*}(5+24C^{\prime}_{F}T_{0})$ gives the first estimate in the lemma for any $\varepsilon\in(0,\varepsilon_{0}]$.
@@
where we have used $\sqrt{x_{r}y_{r}}-b_{r}^{\varepsilon}(t)=\mathcal{O}(1/\sqrt{|\ln\varepsilon|})$ for any $x,y\in\operatorname{supp}\omega^{\varepsilon}(t)$ and (4.3). Therefore, for $R\geqslant 1$
@@ 4.4. Bootstrap argument and ends of the proof of the main results of this section
$\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x\leqslant\frac{3}{4}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
@@
By Proposition 4.11 applied to $\rho(\varepsilon)=\frac{B}{|\ln\varepsilon|-2\ln|\ln\varepsilon|}$, there exists a set $\widetilde{\Omega}_{i}^{\varepsilon}(t)$ such that
@@
$\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\leqslant\frac{h_{0}}{r^{*}}\rho(\varepsilon)\ln|\ln\varepsilon|=\frac{h_{0}}{r^{*}}\frac{B\ln|\ln\varepsilon|}{|\ln\varepsilon|-2\ln|\ln\varepsilon|}\leqslant\frac{3}{4}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|},$
@@
(4.7) $\frac{Bh_{0}}{r^{*}}=20(5+24C_{F}^{\prime}T_{0})\frac{h_{0}}{r*}\leqslant\frac{1}{2}.$
The inequality (4.7) entails the constraint on $h_{0}$ and $T_{0}$ in the main results stated at the beginning of this section. Assuming for instance $T_{0}\leqslant 1/(24C_{F}^{\prime})$ and $h_{0}\leqslant r^{*}/240$, all the results of this section read as follows: there exists $\varepsilon_{0}>0$ depending only on $h_{0}$ such that for any $\varepsilon\in(0,\varepsilon_{0}]$ and every $i\in\{1,\ldots,N\}$, all the inequalities stated hold true for all $t\in[0,T_{\varepsilon}]$. In particular, Proposition 4.1 and respectively Lemma 4.2 follow directly from Proposition 4.11 and resp. Lemma 4.4. Finally, Corollary 4.3 is a direct consequence of Lemma 4.4 combined with the definition of $\Omega_{i}^{\varepsilon}$ (4.2), see Remark 4.5.
@@
The constant $h_{0}$ was chosen small enough so that some constants are absorbed, see (4.7), which explained why we have tracked in all this section the dependence on $h_{0}$. The possibility to choose $h_{0}$ small will not be used below, therefore, $h_{0}$ is fixed as $h_{0}=\min(r^{*}/240,1/2)$ and we will not specify the dependence of the constants on $h_{0}$ in the sequel.
@@
which means that most of the mass can be found in an area of radius $\varepsilon^{1-\eta}$. However, this comes at the price of a relatively rough information on the location of $\hat{\Omega}^{\varepsilon}_{i}(t)$. Specifically, with the localization estimates at hand and the mass being located away from $\hat{\Omega}^{\varepsilon}_{i}(t)$ of order $\mathcal{O}\left(\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)$, we can only show that
@@
by adapting the proof of Lemma 4.4 and no better than that. Therefore, the center of mass can be quite far from the actual location of the vortex filament compared to the size of filament’s core. This is due to the possible filamentation of the vortex filament as discussed in the introduction.
@@ 5. Equation of the motion
Lemma 4.9 provides the equation of the motion of the radial component of $b^{\varepsilon}_{i}(t)$ that holds up to $T_{\varepsilon}$ as $\widetilde{T}_{\varepsilon}=T_{\varepsilon}$ being defined in (4.2), (3.6) respectively. The goal of this section is to derive the full equation for both components and then establish the link with $(X_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$ as in (1.3).
@@ 5.1. Motion of the center of mass
We prove in the following lemma that the motion of the center of mass of the vortex ring is governed by the binormal flow together with the influence of the other vortex rings.
@@
Provided $T_{0}\leqslant 1/(24C^{\prime}_{F})$, there exists $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$\left|\frac{\mathrm{d}}{\mathrm{d}t}b_{i}^{\varepsilon}(t)-\frac{1}{|\ln\varepsilon|}F_{i}^{\varepsilon}(t,b_{i}^{\varepsilon}(t))-\gamma\frac{e_{z}}{4\pi b_{i,r}^{\varepsilon}(t)}\right|\leqslant C\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
@@
As in the beginning of the proof of Lemma 4.9, we start by writing
@@
$+\frac{1}{\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}u_{L}^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$+\frac{1}{\gamma|\ln\varepsilon|}\int u_{R}^{\varepsilon}(t,x)x_{r}\omega(t,x)\mathrm{d}x$
$+\frac{1}{\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}F^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
::=:= $A_{1}+A_{2}+A_{3}+A_{4},$
@@
$A_{4}=\frac{1}{|\ln\varepsilon|}F^{\varepsilon}(t,b^{\varepsilon}(t))+\frac{1}{\gamma|\ln\varepsilon|}\int\big{(}F^{\varepsilon}(t,x)-F^{\varepsilon}(t,b^{\varepsilon}(t))\big{)}x_{r}\omega(t,x)\mathrm{d}x$
then using from Lemma 3.1 that $F^{\varepsilon}$ is Lipschitz and Lemma 4.2
@@
$-\frac{e_{z}}{2\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}\left(\frac{1}{x_{r}^{2}}-\frac{1}{|b_{r}^{\varepsilon}(t)|^{2}}\right)\psi^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x.$
@@
then, by using Corollary 4.10, we obtain that
$\left|A_{2}-\gamma\frac{e_{z}}{4\pi b_{r}^{\varepsilon}(t)}\right|\leqslant C(1+T_{0})\frac{\ln|\ln\varepsilon|}{\ln\varepsilon}\leqslant C\frac{\ln|\ln\varepsilon|}{\ln\varepsilon},$
provided $T_{0}\leqslant 1/(24C^{\prime}_{F})$. ∎
Finally, we show that the velocity $F_{i}^{\varepsilon}$ created by the other vortices behaves as the interaction in the classical point vortex system (see [44, 45]).
@@
Provided $T_{0}\leqslant 1/(24C^{\prime}_{F})$, there exists $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have
@@
$-\frac{e_{z}}{4\pi b_{i,r}^{\varepsilon}(t)}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\sqrt{\frac{y_{r}}{b_{i,r}^{\varepsilon}(t)}}\ln|b_{i}^{\varepsilon}(t)-y|y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y$
$+\frac{1}{b_{i,r}^{\varepsilon}(t)}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\nabla^{\perp}_{x}S(b_{i}^{\varepsilon}(t),y)y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y$
::=:= $A_{1}+A_{2}+A_{3}.$
@@
5.2. Limiting trajectories
We prove in this section that the center of vorticities $(b_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$, defined in (1.6), are located close to $(X_{1\leqslant i\leqslant N}^{\varepsilon})_{i}$ and $(\widetilde{X}_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$, defined in (1.3) and (1.9) respectively.
First, we prove that $(X_{i}^{\varepsilon})_{i}$ and $(\widetilde{X}_{i}^{\varepsilon})_{i}$ are located in $\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}(d_{0},h_{0})$, see (3.4) for the definition of this set, and remain suitably close.